# Polish Medical Exams: A new dataset for cross-lingual medical knowledge transfer assessment

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in handling specialized tasks, including medical problem-solving. However, most studies predominantly focus on English-language contexts. This study introduces a novel benchmark dataset based on Polish medical licensing and specialization exams (LEK, LDEK, PES) taken by medical doctor candidates and practicing doctors pursuing specialization. The dataset was web-scraped from publicly available resources provided by the Medical Examination Center and the Chief Medical Chamber. It comprises over 24,000 exam questions, including a subset of parallel Polish-English corpora, where the English portion was professionally translated by the examination center for foreign candidates. By creating a structured benchmark from these existing exam questions, we systematically evaluate state-of-the-art LLMs, including general-purpose, domain-specific, and Polish-specific models, and compare their performance against human medical students. Our analysis reveals that while models like GPT-4o achieve near-human performance, significant challenges persist in cross-lingual translation and domain-specific understanding. These findings underscore disparities in model performance across languages and medical specialties, highlighting the limitations and ethical considerations of deploying LLMs in clinical practice.
@@ 1 Introduction
The potential of Artificial Intelligence, particularly Large Language Models, is vast, yet these technologies come with considerable risks. One of the most pressing concerns is the phenomenon of “hallucinations”, where LLMs generate responses that appear accurate but are incorrect or misleading. This issue is particularly critical in fields like medicine, where errors can have serious, sometimes life-threatening, consequences. Thus, rigorous evaluation of LLM performance is imperative before their integration into clinical practice Minaee et al. (2024).
@@
LLM performance varies widely due to differences in training methods, datasets, and objectives. These factors ultimately dictate how well an LLM can execute specific tasks. The quality and diversity of training datasets are especially crucial for LLM performance in specialized domains, such as medicine Minaee et al. (2024). Models trained on comprehensive, domain-specific datasets tend to outperform those trained on general-purpose data, although this view has been questioned by many.
@@
Language also plays a significant role in the performance of LLMs. Many of the most widely studied models are trained using multilingual datasets that predominantly contain English texts. Consequently, LLMs often exhibit better performance with English-language inputs, while struggling with non-English content Minaee et al. (2024). Moreover, LLMs trained solely on non-English texts may miss crucial knowledge available only in English.
Modern medicine is fundamentally evidence-based, and one might presume that the correct management of a certain medical problem should be nearly universal worldwide. However, in practice, the situation is considerably more complex. Clinical practices are influenced by a complex interplay of factors, including epidemiological trends, economic resources, healthcare infrastructure, technology availability, cultural norms, legal and regulatory frameworks, and even environmental conditions Minaee et al. (2024).
Epidemiology shapes medical education, with curricula typically emphasizing conditions prevalent in the local population. Medical students are trained to manage health problems that are common in their geographical region. For example, in areas with high tuberculosis prevalence, medical education places significant focus on its diagnosis and management. Conversely, regions facing a higher prevalence of non-communicable diseases tend to emphasize training in chronic, lifestyle conditions. This localized focus ensures healthcare professionals are adequately prepared for the most relevant health issues within their communities. However, it can also mean that language-specific training data, when applied to other populations, may under-represent certain conditions.
Furthermore, physicians often rely on the local prevalence and incidence of diseases to guide their diagnostic decisions. For instance, when the clinical presentation is non-specific, in regions where infectious diseases are common, healthcare professionals are more likely to suspect an infectious cause, whereas in developed regions, noninfectious etiologies may be considered earlier.
Treatment choices are similarly influenced by local health data, for example when choosing antibiotics without specific microbiological test results (an approach called empiric antibiotic therapy). Doctors often decide on antibiotic combinations based on how resistant local bacteria are to these drugs. For example, the treatment to eliminate Helicobacter pylori, a bacteria which can cause stomach ulcers, differs depending on regional resistance to common antibiotics like clarithromycin and metronidazole. In areas where clarithromycin is less effective, doctors may use a more complex combination of drugs to increase success rates Yang et al. (2014). These differences show why treatment plans need to be adapted to local conditions for the best patient outcomes.
Moreover, the availability of diagnostic methods and medications varies across different regions of the world, leading to differences in clinical practice standards. For instance, if a blood test result justifies initiating a specific treatment, in regions with widespread access to laboratory testing, the standard approach would typically involve performing the test first and basing the treatment decision on its result. Conversely, in countries with limited access to laboratory diagnostics, the standard practice might involve initiating treatment without prior testing and discontinuing it if no improvement is observed. Clinical decision-making often hinges on weighing potential benefits against risks. In the example above, initiating therapy without laboratory confirmation may not necessarily be considered a mistake, especially if the medication is safe and testing is significantly limited.
Societal norms and legal frameworks significantly influence medical practice, sometimes diverging from current medical evidence. In regions with restrictive legislation, access to procedures such as abortion or gender-affirming surgeries may be heavily restricted. Conversely, certain procedures like elective cosmetic surgeries are widely practiced in some countries. As a result, the corpora used to train LLMs can embed cultural and legal influences, potentially skewing the model’s recommendations towards decisions aligned with the dominant perspectives in the data. Additionally, medical examinations can also be biased, as these norms and laws can influence the framing of questions and the designation of correct answers.
@@
LLMs trained predominantly on English-language data are more likely to align with the disease prevalence and clinical guidelines typical of English-speaking countries. As a result, their diagnostic and therapeutic recommendations may be biased towards practices common in these regions. On the other hand, when faced with the same clinical scenario presented in different languages, an LLM may yield dissimilar responses, reflecting the variety of healthcare practices across multiple countries that appear in the training data. This phenomenon becomes especially pertinent when evaluating LLM performance on non-English medical tests, such as those conducted in Poland, where the disease prevalence and medical guidelines may diverge from those in English-speaking countries.
@@
To study LLM performance in the medical setting, we propose a new benchmark that consists of publicly available exam questions from medical and dental licensing examinations, as well as specialist-level exams conducted in Poland. [^1] The dataset is available at https://huggingface.co/spaces/amu-cai/Polish_Medical_Exams. The dataset comprises over 24,000 questions, primarily in Polish. A subset of questions from the licensing examinations also includes English-language counterparts, facilitating comparative analysis.
We use the proposed benchmark to study the behavior of LLMs aiming at the following research questions:
@@
  • How does the performance of evaluated LLMs compare to that of human doctors and medical students?
  • To what extent do LLMs provide divergent responses to general medical examination questions in Polish versus English, considering high-quality human experts’ translations?
  • What are the differences in the performance of LLMs on general versus specialized Polish medical examinations?
  • How effectively do LLMs handle examination questions across various medical specialties (e.g., cardiology, neurology)?
@@ 2 Related work
LLMs are poised to transform various aspects of medicine by supporting medical professionals and enhancing research and education Abd-Alrazaq et al. (2023); Clusmann et al. (2023). They can assist with literature summarization, data extraction, manuscript drafting, patient-clinical trial matching, and the creation of educational content Clusmann et al. (2023); Harrer (2023); Yang et al. (2023). By facilitating the conversion of unstructured to structured data, they streamline communication, translating complex medical information and summarizing patient records to simplify documentation Clusmann et al. (2023). Applications in medical report generation and preauthorization letters can notably reduce the administrative burden on clinicians, allowing more focus on patient care Harrer (2023). This enhanced efficiency benefits healthcare systems not only economically but also supports personalized, patient-centered care through improved clinician-patient interactions Clusmann et al. (2023); Nazi and Peng (2024). Additionally, LLMs may aid diagnostics and management planning by analyzing large volumes of medical data and monitoring patient parameters Nazi and Peng (2024).
The integration of LLMs in healthcare necessitates rigorous evaluation to ensure reliability and safety, given the complexity and high stakes of medical decisions. Robust evaluation is crucial for assessing the performance of LLMs, identifying weaknesses, and mitigating biases to maintain patient safety and promote equity Karabacak and Margetis (2023); Li et al. (2023). Further, evaluation should ensure that LLMs genuinely enhance clinical care and effectively support healthcare professionals Karabacak and Margetis (2023).
Evaluating LLMs for medical use requires a tailored approach that goes beyond standard metrics, incorporating the specific demands of healthcare. Evaluation efforts should consider both technical performance, such as accuracy, reasoning abilities, and factual reliability using benchmark datasets (e.g., medical licensing exams), and real-world utility, including clinical impact studies and workflow integration Karabacak and Margetis (2023); Chang et al. (2024); Nazi and Peng (2024).
Medical datasets are created using various approaches. The MedQA dataset comprises United States Medical Licensing Examination (USMLE)-style questions Jin et al. (2020), while JAMA Clinical Challenge is based on the JAMA Network Clinical Challenge archive Chen et al. (2024). Medbullets uses simulated clinical questions sourced from social media posts Chen et al. (2024), and PubMedQA utilizes questions and contexts derived from PubMed articles Jin et al. (2019).
@@
For Polish, Lekarski Egzamin Końcowy (LEK, Eng. Medical Final Examination) was used as a benchmark Rosoł et al. (2023); Bean et al. (2024); Suwała et al. (2023). LEK is available in both Polish and English, allowing researchers to evaluate the influence of language on LLM performance. To date, analyses have primarily focused on GPT models, though several other LLMs, including LLaMa and Med42, have also been evaluated Bean et al. (2024).
Regarding the Państwowy Egzamin Specjalizacyjny (PES, Eng. State Specialisation Examination), a few studies have assessed GPT’s performance in specialized field exams Suwała et al. (2023); Kufel et al. (2023); Wojcik et al. (2023). Pokrywka et al. (2024) provided a comprehensive evaluation of GPT-3.5 and GPT-4 on the PES, utilizing 297 exams across 57 specialties.
Jin et al. (2024) proposed a benchmark for the cross-lingual evaluation of LLMs. However, the questions they used were translated by a machine translation system, while the questions in our benchmark were translated by human medical experts. Furthermore, we evaluated new models that demonstrate much better performance Kipp (2024).
3 Polish Medical Exams
In Poland, five types of exams for physicians and dentists are conducted: LEK (Lekarski Egzamin Końcowy, Eng. Medical Final Examination), LDEK (Lekarsko-Dentystyczny Egzamin Końcowy, Eng. Dental Final Examination), LEW (Lekarski Egzamin Weryfikacyjny, Eng. Medical Verification Examination), LDEW (Lekarsko-Dentystyczny Egzamin Weryfikacyjny, Eng. Dental Verification Examination), and PES (Państwowy Egzamin Specjalizacyjny, Eng. National Specialization Examination, Board Certification Exam). LEW and LDEW are for graduates of medical or dental studies carried outside of the European Union. Passing these exams is necessary for them to legally practice in Poland. [^2] https://www.cem.edu.pl/lew_info.php
@@ 3 Polish Medical Exams
However, these LEW and LDEW are taken by a relatively small number of candidates, and access to previous exam questions is limited. Therefore, they are not included in our work.
https://www.cem.edu.pl/ldek_info.php
Both LEK and LDEK are four-hour exams conducted twice a year. Each exam consists of 200 multiple-choice questions with five possible answers, of which only one is correct. The questions cover a wide range of medical or dental disciplines. The distribution of questions from various fields is presented in Tables 1 and 2. To pass, a candidate must correctly answer at least 56% of the questions. Physicians and dentists can retake these exams multiple times, even after passing, if they are dissatisfied with their score. [^4] https://www.cem.edu.pl/lek_info.php
https://www.cem.edu.pl/ldek_info.php
A controversial rule has been introduced in 2022, stipulating that 70% of the exam questions come from a publicly available database, which includes 2,870 questions for LEK and 3,198 for LDEK. After these changes, the average exam scores and the percentage of passing candidates increased significantly. [^5] https://www.cem.edu.pl/lep_s_h.php
https://www.cem.edu.pl/ldep_s_h.php
Discipline Questions
Internal medicine* 39
Pediatry* 29
Surgery* 27
Obstetrics and gynecology* 26
Psychiatry 14
Family medicine* 20
Emergency medicine and intensive care 20
Bioethics and medical law 10
Medical certification 7
Public health 8
Discipline Questions
Conservative dentistry* 46
Pediatric dentistry* 29
Oral surgery* 25
Prosthetic dentistry 25
Periodontology* 20
Orthodontics* 20
Emergency medicine 10
Bioethics and medical law 10
Medical certification 7
Public health 8
4 Dataset overview
The dataset comprises medical exams from the Medical Examination Center (CEM) and the Supreme Medical Chamber (NIL), covering LEK, LDEK, and PES exams from 2008–2024. The exams were sourced as HTML quizzes and PDF files, with missing data from 2016–2020 (LEK/LDEK) and 2018–2022 (PES) partially filled using archives published on the NIL website. The exams were categorized by specialization, with questions and answers stored separately. Automated tools were used to scrap and process data, balancing parallelization with server constraints. Preprocessing ensured the dataset’s suitability for text-only AI benchmarks by removing irrelevant files, questions containing images, and content misaligned with current medical knowledge. We refer to these as "invalidated questions" throughout the text. Detailed descriptions of data sources, acquisition methods, and quality considerations are provided in Appendix B.
@@ 4 Dataset overview
Finally, we created five sub-datasets: LEK, LDEK, PES, LEK en (LEK translated into English), and LDEK en (LDEK translated into English). Not all of them were released in the same edition, particularly the Polish and English counterparts. Therefore, the results presented in Section 5 should not be used to directly compare LLM performance on Polish exams with their English translations. To address this, we focused on the overlapping years and reported these results in Section 7. For PES dataset, we collected a total of 180,712 questions. For our analysis, we selected only the latest exam from each specialty and based our analysis on these. Detailed dataset statistics are provided in Table 3. In summary, our analysis encompasses over 24,000 questions. For LLM inference, we utilized the Huggingface Transformers library Wolf (2019) and the OpenAI API.
@@ 5 Performance of LLMs on exams
We categorized the models under study into the following groups: medical LLMs (models fine-tuned on English medical data), general-purpose multilingual LLMs, Polish-specific models, and models with restricted APIs.
@@
General-Purpose Multilingual Models: Qwen2.5 Instruct (7B and 72B versions) Team (2024), Llama-3.1 Instruct (8B and 70B versions), Llama-3.2-3B Dubey et al. (2024), mistralai/Mistral-Small-Instruct-2409, and Mistral-Large-Instruct-2407 Jiang et al. (2023).
@@
LLMs were evaluated by directly prompting them to answer exam questions. Each prompt included a brief introduction stating that the task was an exam for medical professionals consisting of single-choice questions. No additional examples or explanations were provided in the prompt; specifically, few-shot prompting was not employed. We believe this approach is appropriate for evaluating the models in a setting closely resembling the actual human exam environment.
We report the models’ results as the percentage of correct answers in Table 4 and the number of exams passed in Table 5. Our findings are as follows: GPT-4o is the best performing model overall. Particularly in the PES category, it outperforms the second-best model. GPT-4o is capable of passing all evaluated exams except for six PES exams. However, GPT-4o-mini performs significantly worse than GPT-4o and is also inferior to general-purpose open models. Among the non-restricted API models, Meta-Llama-3.1-70B-Instruct is the best performer. Generally, general-purpose models outperform medical-specific models, possibly because the latter were fine-tuned on English medical data. The Polish-specific general-purpose model, Bielik-11B-v2.2-Instruct, performs worse than the top multilingual general-purpose models such as Meta-Llama-3.1-70B-Instruct, Qwen2.5-72B-Instruct, and Mistral-Large-Instruct-2407. However, for scenarios where deployment costs are more critical than performance, Bielik-11B-v2.2-Instruct may be preferable, as it still outperforms Meta-Llama-3.1-8B-Instruct of similar size in Polish-only exams. Our final recommendation is to use GPT-4o for Polish medical data tasks. If using a restricted API is not feasible (e.g., due to patient anonymity requirements), Meta-Llama-3.1-70B-Instruct is suggested as an alternative.
@@
6 Specialty performance on PES exams
Among the 72 unique PES specialties, certain areas of medicine consistently challenge the majority of tested models, while others frequently rank among the highest-scoring categories based on model accuracy. By identifying the top five highest and lowest-scored categories, we gain insights into specific domains where models excel or struggle, highlighting their potential limitations in these fields. Their detailed results are presented in Appendix A.

### v2

# Polish-English medical knowledge transfer: A new benchmark and results

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in specialized tasks, including medical problem-solving. However, most studies predominantly focus on English-language contexts. This study introduces a novel benchmark dataset based on Polish medical licensing and specialization exams (LEK, LDEK, PES). The dataset, sourced from publicly available materials provided by the Medical Examination Center and the Chief Medical Chamber, includes Polish medical exam questions, along with a subset of parallel Polish-English corpora professionally translated for foreign candidates. By structuring a benchmark from these exam questions, we evaluate state-of-the-art LLMs, spanning general-purpose, domain-specific, and Polish-specific models, and compare their performance with that of human medical students and doctors. Our analysis shows that while models like GPT-4o achieve near-human performance, challenges persist in cross-lingual translation and domain-specific understanding. These findings highlight disparities in model performance across languages and medical specialties, emphasizing the limitations and ethical considerations of deploying LLMs in clinical practice.
@@ 1 Introduction
The potential of Artificial Intelligence, especially Large Language Models (LLMs), is vast, but they come with considerable risks, particularly the issue of “hallucinations”, where LLMs produce incorrect or misleading responses. This is especially concerning in fields like medicine, where errors can have serious consequences. Therefore, rigorous evaluation of LLM performance is essential before their clinical integration Minaee et al. (2024).
@@
LLM performance varies significantly due to differences in training methods, datasets, and objectives, which affect their ability to perform specific tasks. The quality and diversity of training datasets are particularly important for specialized domains like medicine Minaee et al. (2024). While models trained on comprehensive, domain-specific datasets are expected to outperform those trained on general-purpose data, this assumption has been challenged Nori et al. (2023).
@@
Language also significantly impacts LLM performance. Most widely studied models are trained on multilingual datasets, predominantly in English, leading to better performance with English-language inputs and challenges with non-English content Minaee et al. (2024). Additionally, LLMs trained exclusively on non-English texts may lack important knowledge available only in English.
Modern medicine is evidence-based, and one might assume that the correct management of medical issues should be nearly universal. However, in practice, clinical practices are shaped by various factors, leading to significant variations in medical guidelines across countries. For instance, Zhou et al. (2024) analyzed 22 clinical practice guidelines from 15 countries, highlighting notable differences in recommendations for managing lower back pain.
@@
LLMs trained primarily on English-language data are likely to align with disease prevalence and clinical guidelines typical of English-speaking countries. Consequently, their diagnostic and therapeutic recommendations may be biased towards practices common in these regions. When presented with the same clinical scenario in different languages, an LLM may produce varying responses, reflecting the diversity of healthcare practices across countries represented in the training data. Such discrepancies could be revealed by evaluating LLMs on non-English medical tests, like those conducted in Poland, where disease prevalence and medical guidelines may differ from those in English-speaking countries.
@@
To primarily assess the performance of LLM models in medical question-answering tasks, we introduce a new benchmark based on publicly available exam questions from medical and dental licensing exams, as well as specialist-level exams conducted in Poland [^1] The dataset is available at https://huggingface.co/spaces/amu-cai/Polish_Medical_Exams. This dataset includes over 22,000 questions, primarily in Polish, with a subset of licensing exam questions also available in English, enabling comparative analysis. We propose a benchmark that enables the study of LLM behavior by addressing the following research questions:
@@
  • How do LLMs compare to human doctors and medical students in performance?
  • How do LLMs’ responses differ to general medical questions in Polish versus English, based on high-quality expert translations?
  • What are the differences in the performance of LLMs on general versus specialized Polish medical exams?
  • How well do LLMs handle questions across various medical specialties (e.g., cardiology, neurology)?
@@ 2 Related work
LLMs have the potential to revolutionize medicine by assisting medical professionals in key areas such as medical education, literature summarization, data extraction, manuscript drafting, and patient-clinical trial matching Harrer (2023); Yang et al. (2023). They streamline communication by converting unstructured data to structured formats and simplifying documentation, such as summarizing patient records and generating medical reports Clusmann et al. (2023). This reduces administrative burdens on clinicians, allowing more focus on patient care Harrer (2023). LLMs also enhance personalized, patient-centered care, improve clinician-patient interactions, and may aid in diagnostics and management planning by analyzing medical data and monitoring patient parameters Clusmann et al. (2023); Nazi and Peng (2024).
Integrating LLMs into healthcare requires thorough evaluation to ensure reliability, safety, and equity, while identifying weaknesses and addressing biases to improve clinical care and support healthcare professionals Karabacak and Margetis (2023); Li et al. (2023). This evaluation should go beyond traditional performance metrics to include factors such as accuracy, reasoning, and factual reliability, using benchmarks like medical licensing exams, as well as assessing real-world utility, including clinical impact and workflow integration Chang et al. (2024).
Small, fine-tuned BERT-style models continue to outperform LLMs in certain NLP tasks, such as text classification Bucher and Martini (2024). However, the emergence of LLMs, such as GPT-3.5 and Med-PaLM 2, has led to significant advancements in medical question-answering benchmarks, including MedQA Jin et al. (2021), MedMCQA Pal et al. (2022), and PubMedQA Jin et al. (2019). For this specific NLP task, general-purpose LLMs enhanced with specialized prompting strategies Nori et al. (2023) or fine-tuned domain-specific models surpass small encoder-only models in performance Singhal et al. (2025).
@@
For Polish, Lekarski Egzamin Końcowy (LEK, Eng. Medical Final Examination) is used as a benchmark Rosoł et al. (2023); Bean et al. (2024); Suwała et al. (2023). LEK is available in both Polish and English, allowing researchers to evaluate the influence of language on LLM performance. To date, analyses have primarily focused on GPT models, though several other LLMs, including LLaMa and Med42, have also been evaluated Bean et al. (2024).
Regarding the Państwowy Egzamin Specjalizacyjny (PES, Eng. Polish Board Certification Examination), a few studies have assessed GPT’s performance in specialized field exams Suwała et al. (2023); Kufel et al. (2023); Wojcik et al. (2023). Pokrywka et al. (2024) provided a comprehensive evaluation of GPT-3.5 and GPT-4 on the PES, utilizing 297 exams across 57 specialties in Polish. We extend this research by incorporating additional PES specialties and introducing new exam types—the LEK and LDEK—in both Polish and English. Our study also includes cross-lingual evaluation of LLMs, comparisons with human performance, and assessments of publicly available LLMs, which were not done by Pokrywka et al. (2024).
Jin et al. (2024) proposed a benchmark for the cross-lingual evaluation of LLMs. However, the questions they used were translated by a machine translation system, while the questions in our benchmark are translated by human medical experts from the examination center. Furthermore, we evaluated new models that demonstrate much better performance Kipp (2024).
3 Polish medical exams dataset overview
The LEK (Lekarski Egzamin Końcowy, Medical Final Examination) is a standardized exam for medical graduates and final-year students in Poland. Passing this exam, along with completing a postgraduate internship, is mandatory to obtain a medical license. Starting from 2022, 70% of the questions come from a publicly available database, which includes 2,870 questions for LEK. The exam is conducted twice a year and lasts four hours, consisting of 200 multiple-choice questions. Candidates are allowed to retake the exam multiple times, even after passing, to improve their scores.
The LDEK (Lekarsko-Dentystyczny Egzamin Końcowy, Dental Final Examination) is the equivalent exam for dentistry graduates and final-year students, following the same format and requirements as the LEK.
The PES (Państwowy Egzamin Specjalizacyjny, National Specialization Examination) is a mandatory exam for physicians and dentists who have completed specialization training, including required internships and courses. It consists of a written test and an oral examination. The written test, held twice a year for each specialty, typically includes 120 multiple-choice questions, with one correct answer per question, and a passing score of 60%. Candidates achieving at least 70% on the written part are exempt from the oral examination, a rule introduced in late 2022. PES is considered the most challenging exam in the professional career of a medical doctor in Poland, and unlike LEK and LDEK, its questions are not made public before the exam.
In Poland, five types of exams for physicians and dentists are conducted: LEK (Lekarski Egzamin Końcowy, Eng. Medical Final Examination), LDEK (Lekarsko-Dentystyczny Egzamin Końcowy, Eng. Medical-Dental Final Examination), LEW (Lekarski Egzamin Weryfikacyjny, Eng. Medical Verification Examination), LDEW (Lekarsko-Dentystyczny Egzamin Weryfikacyjny, Eng. Medical-Dental Verification Examination), and PES (Państwowy Egzamin Specjalizacyjny, Eng. National Specialization Examination, Board Certification Exam). LEW and LDEW are for graduates of medical or dental studies carried outside of the European Union. Passing these exams is necessary for them to legally practice in Poland. [^2] https://www.cem.edu.pl/lew_info.php
@@ 3 Polish medical exams dataset overview
However, these LEW and LDEW are taken by a relatively small number of candidates, and access to previous exam questions is limited. Therefore, they are not included in our work. The extensive descriptions of medical exams are included in Appendix A.
The dataset comprises medical exams from the Medical Examination Center (Centrum Egzaminów Medycznych - CEM) and the Supreme Medical Chamber (Naczelna Izba Lekarska - NIL), covering LEK, LDEK, and PES exams from 2008–2024. It sources the exams as HTML quizzes and PDF files, with missing data from 2016–2020 (LEK/LDEK) and 2018–2022 (PES) partially filled using archives published on the NIL website. The exams are categorized by specialization, with questions and answers stored separately. Automated tools scrape and process the data, balancing parallelization with server constraints. Preprocessing ensures the dataset’s suitability for text-only AI benchmarks by removing irrelevant files, questions containing images, and content misaligned with current medical knowledge. We refer to these as "invalidated questions" throughout the text. Detailed descriptions of data sources, acquisition methods, and quality considerations appear in Appendix E.
@@
Finally, we create five sub-datasets: LEK, LDEK, PES, LEK en (LEK translated into English), and LDEK en (LDEK translated into English). Not all of them are released in the same edition, particularly the Polish and English counterparts. Therefore, the results presented in Section 4 should not be used to directly compare LLM performance on Polish exams with their English translations. To address this, we focus on the overlapping years and report these results in Section 5. For the PES dataset, we collected a total of 180,712 questions. For the analysis in Sections 4, 5, and 6, we select only the most recent exam from each specialty and base our analysis on these exams. Detailed dataset statistics are provided in Table 1, and example questions are presented in Appendix B. In total, our analysis covers over 22,000 questions. For LLM inference, we use the Huggingface Transformers library Wolf (2019) and the OpenAI API.
@@ 4 Performance of LLMs on exams
We categorize the models under study into the following groups: medical LLMs (models fine-tuned on English medical data), general-purpose multilingual LLMs, Polish-specific models, and models with restricted APIs.
@@
General-Purpose Multilingual Models: Qwen2.5 Instruct (7B and 72B versions) Team (2024), Llama-3.1 Instruct (8B and 70B versions), Llama-3.2-3B Instruct Dubey et al. (2024), mistralai/Mistral-Small-Instruct-2409, and Mistral-Large-Instruct-2407 Jiang et al. (2023).
@@
We evaluate LLMs by directly prompting them to answer exam questions. Each prompt includes a brief introduction stating that the task is an exam for medical professionals consisting of single-choice questions. We do not provide additional examples or explanations in the prompt, and we do not use few-shot prompting. This approach aligns with the actual human exam environment, making it suitable for evaluating the models. Check C for the exact prompts in Polish and English.
We report the models’ results as the percentage of correct answers in Table 2 and the number of exams passed in Table 3. Our findings are as follows: GPT-4o is the best performing model overall. Particularly in the PES category, GPT-4o outperforms the second-best model, Meta-Llama-3.1-70B-Instruct. GPT-4o is capable of passing all evaluated exams except for six PES exams. However, GPT-4o-mini performs significantly worse than GPT-4o and is also inferior to general-purpose open models. Among the open source models, Meta-Llama-3.1-70B-Instruct is the best performer. General-purpose models outperform medical-specific models, possibly because the latter were fine-tuned on English medical data. The Polish-specific general-purpose model, Bielik-11B-v2.2-Instruct, performs worse than the top multilingual general-purpose models such as Meta-Llama-3.1-70B-Instruct, Qwen2.5-72B-Instruct, and Mistral-Large-Instruct-2407. However, for scenarios where deployment costs are more critical than performance, Bielik-11B-v2.2-Instruct may be preferable, as it still outperforms Meta-Llama-3.1-8B-Instruct of similar size in Polish-only exams. Our final recommendation is to use GPT-4o for Polish medical data tasks. If using a restricted API is not feasible (e.g., due to patient anonymity requirements), Meta-Llama-3.1-70B-Instruct is suggested as an alternative.
@@
The performance of LLMs varies significantly based on specialization in PES exams, which was noted by Pokrywka et al. (2024) before. We provide a detailed analysis across specialties in Appendix D, expanding upon the previous authors’ findings with LLM other than the GPT family.